# HARMONIC: Cognitive and Control Collaboration in Human-Robotic Teams

### v1

@@ VI-A High-level planning
The ontologies of agents in HARMONIC contain scripts for complex events that a particular agent can perform or understand other agents performing. Plans are as instances of scripts, with parameter values set for the situation at hand. Plans are instantiated by executing one or more of ontologically stored meta-plans, that is, plans that HARMONIC robots instantiate to guide the generation of plans in their operational domain. One such meta-plan is the collaborative activity plan, whose purpose is to help agents operating in teams organize themselves to accomplish a goal.
@@
The details of the collaborative activity plan vary based on the agent’s role in the team. For example, the team (or task) leader’s plan is typically more complex than the plans of its subordinates since team leaders select the plan to accomplish the target goal, resolve preconditions, and instruct their subordinates. Subordinates’ plans, by contrast, include awaiting instructions from the leader and, passively absorbing any other related information that comes their way.
@@
The leader’s script for collaboration is shown below. Since this is a meta-plan, it can be applied to many specific plans and, starts with few details populated. Part of its task is to fill in task-specific details. Select comments are provided for this human-readable formalism. For further information, please refer to [34, 35, 7].
@@
As the leader begins to process the meta-plan, it will select the domain plan to pursue. In the example below, the selected plan is SEARCH-FOR-LOST-OBJECT. The details of the selected plan can be placed into the [RUN-PLAN] section of collaborative activity, and any preconditions that must be resolved can be added to the [PRECONDITIONS] section. An excerpt from the leader’s script for COLLABORATIVE-ACTIVITY is shown below.
@@
VII Evaluation
VII-A Simulation Environment and Task Design
We evaluated the HARMONIC framework’s teaming strategy using a simulated search and retrieval task set in an apartment environment, as illustrated in Figure 5. The scenario involves a HMR team consisting of an unmanned ground vehicle (UGV), a drone, and a human team member serving as the team leader. The objective is to find a set of lost keys. The human initiates the task by remotely communicating with the robotic team members.
The simulated apartment contains zones with varied robot accessibility, necessitating coordinated efforts. We categorize these as: a) spaces accessible to both robots, b) elevated surfaces accessible to only the drones, and c) spaces below surfaces accessible only to the UGVs, as shown in Fig. 3. Each zone comprises waypoints - strategically defined locations that are dynamically utilized by the strategic layer. For this evaluation, we manually designated these waypoints and zones. However, in the future this process will be automated using machine learning models to classify and map zones based on spatial characteristics and accessibility, which will enhance the system’s adaptability to new environments.
Figure 3: Zones in the search environment
The simulation environment, developed using the Unity game engine, incorporates abstracted versions of robot behaviors to reduce complexity. This approach allowed us to focus on key aspects of the HARMONIC framework’s performance while simplifying the implementation of complex robot actions that are not central to our current work.
VII-B Heterogenous Robot Team Configuration
The simulations involved two robots, an Unmanned Ground Vehicle (UGV) and a drone (Fig. 4). These robots are both functionally and structurally heterogeneous [44], thus providing a diverse range of complementary perspectives and capabilities to the team. Both robots are equipped with a suite of sensors that enable them with localization, collision detection, and object recognition capabilities.
@@ VII-B Heterogenous Robot Team Configuration
Figure 4: (a) UGV for ground-level exploration. (b) Parrot drone conducting an aerial scan of objects.
Each robot runs its own instances of the tactical and strategic modules, enabling decentralized planning and control within the team. For the current scenario, the UGV is the robot team leader at the task level, making it responsible for maintaining interactions with the human and the drone. Currently, the leader assignment is randomized. In future work, we intend to develop and evaluate heuristic methods for implementing dynamic team hierarchies. This approach will optimize leadership selection based on contextual factors and individual agent capabilities depending on the task.
In the tactical layer of both robots, sensor data is encoded into data frames and periodically shared with the strategic layer. Additionally, each robot incorporates reactive control strategies for collision avoidance, that are embedded within behavior trees in their respective tactical layers.
The UGV is a custom-designed robot featuring a manipulator mounted atop a multidirectional mobile platform, as illustrated in Fig. 4.a. The robot’s design allows it to access all ground-level locations and maneuver under certain objects, enhancing its capability to perform intricate search and manipulation tasks in confined spaces. Figure 2.c presents an excerpt from the Behavior Tree (BT) running on the Unmanned Ground Vehicle (UGV). This BT is based on the template shown in Figure 2.a.
@@
Figure 5: (1) Chat transcript between Danny and the robots. (2) UGV’s (leader) Text Meaning Representations (TMRs) (3, 6) A sample Vision Meaning Representations (VMRs) of detected objects by the robots (5) Leader’s agenda. (4, 7) Complete thought transcripts of UGV and Drone.
The drone, on the other hand, can effectively perform search operations by scanning through a wide range of environments with sensors to scan the objects below it, as shown in Fig. 4.b. Its vertical mobility allows it to navigate over obstacles and access areas that may be challenging for the UGV (type-b zones).
Since the robots communicate in natural language, their dialog and actions are comprehensible to humans, and humans can intervene when necessary.
VII-C Collaborative Search for Lost Keys
The simulation environment is connected to a chatroom interface through which the human (Danny) communicates with the robots. A user interface (see Fig. 5) includes not only a chat window but additional widgets that display the robots’ interpretations of language (TMRs) and visual inputs(VMRs), natural language renderings of the robots’ thoughts, and their goal and plan agendas. This arrangement illustrates the transparency of the robots’ operations and provides insights into their cognitive processes. The discussion below makes reference to the labels in Fig. 5  [^1] A video of the complete simulation can be found at https://youtu.be/LpuE0yN0SEI
VII-C1 Task Initiation
@@
The scenario begins with Danny, initiating the task by sending the robots a message (M1). This input triggers the team leader (UGV) to place a COLLABORATIVE-ACTIVITY on its agenda (Fig. 5.5) and launch the SEARCH-FOR-LOST-OBJECT plan. Concurrently, the drone awaits instructions.
VII-C2 Information Gathering
@@
The UGV proceeds to verify preconditions for the SEARCH-FOR-LOST-OBJECT plan (Fig. 5.5). Although the object type is already known, information about its features and last-seen location would be useful. The UGV queries Danny for this information (M2/M4), and receives responses (M3/M5). The latter response prompts the system to prioritize the entry-way sub-zone (type-a zone) in the search sequence.
VII-C3 Search Execution
@@
With preconditions met, the UGV instructs the drone to initiate the search (M6), and both robots begin exploring their assigned areas. The individual tactical modules on the robots control the search process using a waypoint strategy, while the strategic (cognitive) module maintains awareness of area existence without directly guiding robot navigation. This approach allows for efficient local path planning while preserving high-level planning in the strategic layer.
VII-C4 Communication and Coordination
@@
Throughout the search, robots report their findings to each other and the human (M7-9). When a robot fails to locate the keys in a searched area, it communicates this to its partner.
VII-C5 Task Completion
@@
The VMRs widget (Fig. 5.3 and Fig. 5.6) displays the object detection results that the strategic layer processes from sensing frames communicated by the tactical layer. During the search execution, the strategic module continuously analyzes the sensor data frames, grounding these VMRs against the instance of the KEY object stored in its episodic memory. When the features match, the search is halted by either of the robots (leader in this case), informs the team (M8), and reports to Danny (M9). Notably, the UGV uses different language constructs when communicating with the drone versus Danny in (M9), demonstrating the cognitive agent’s ability to generate context-appropriate language.
VIII Conclusion and Future Work
@@
This work presents a step towards the development of cognitive robots capable of effectively collaborating in heterogeneous human-robot teams. The HARMONIC architecture, with its dual-layer approach that combines strategic cognitive processing and tactical control enables the embodiment of cognitive architectures enhancing robots to exhibit metacognitive abilities, including team task understanding, episodic memory utilization, and multimodal perception interpretation. By integrating natural language communication and reasoning capabilities, our approach facilitates complex team interactions, enhances explainability, and builds trust the key factors in human-robot collaboration.
@@
The glass-box nature of our system, unlike black-box models based on LLMs or VLAs, provides transparency and interpretability to modules whose operation and output should be explainable to humans. Our demonstration system of a simulated search task involving two robots and a human teammate illustrates the potential for HARMONIC to be applied to real-world applications. Future research will focus on expanding the range of tasks and scenarios in which HARMONIC agents can successfully operate; enhancing the robots’ learning capabilities; conducting real-world trials using physical robots to validate the framework’s effectiveness in diverse operational contexts; developing collaborative online planning strategies; and, enhancing real-time adaptability in dynamic environments.

### v2

@@ IV-B1 High-level planning
The ontologies of agents in HARMONIC contain scripts for complex events that a particular agent can perform or understand other agents performing. Plans are instances of scripts with parameter values set for the situation at hand. These are instantiated by executing one or more of the ontologically stored meta-plans, that is, plans that HARMONIC robots instantiate to guide the generation of plans in their operational domain. One such meta-plan is the COLLABORATIVE-ACTIVITY plan, whose purpose is to help agents operating in teams organize themselves to accomplish a goal.
@@
The details of the COLLABORATIVE-ACTIVITY plan vary based on the agent’s role in the team. For example, the team (or task) leader’s plan is typically more complex than the plans of its subordinates since team leaders select the plan to accomplish the target goal, resolve preconditions, and instruct their subordinates. Subordinates’ plans, by contrast, include awaiting instructions from the leader and passively absorbing any other related information that comes their way.
@@
The leader’s script for collaboration is shown below. Since this is a meta-plan, it can be applied to many specific plans and starts with few details populated. Part of its task is to fill in task-specific details. Select comments are provided for this human-readable formalism. For further details on this formalism, please refer to [11, 12, 9].
@@
As the leader begins to process the meta-plan, it will select the domain plan to pursue. In the example below, the selected plan is SEARCH-FOR-LOST-OBJECT. The details of the selected plan can be placed into the [RUN-PLAN] section of COLLABORATIVE-ACTIVITY, and any preconditions that must be resolved can be added to the [PRECONDITIONS] section. An excerpt from the leader’s script for COLLABORATIVE-ACTIVITY is shown below.
@@
IV-B2 Low-level control
Behavior Trees (BTs) in the tactical layer for both robots follow a standardized design template prioritizing safety and operational needs as shown in Fig. 3.c. This design leverages left-to-right execution priority, ensuring the robot’s critical needs and priorities. Collision avoidance sub-trees are positioned leftmost, followed by robot-needs sub-trees, then action-command sub-trees. On the far right, fallback sub-trees handle behaviors like random exploration or returning to the base station. An excerpt from the UGV’s BT is shown in Fig. 3.d.
@@ IV-B2 Low-level control
Figure 3: (a) UGV for ground-level exploration. (b) Parrot drone conducting aerial scan. (c) BT design template for the tactical layer of robots. (d) Sample BT on UGV.
@@
Figure 4: (1) Chat transcript between Danny (human) and the robots. (2) UGV’s (leader) sample Vision Meaning Representations (VMRs) of detected objects by the robots (3) Leader’s agenda. (4, 5) Complete thought transcripts of UGV and Drone. (6) Multiple instances from the simulated apartment environment showing the UGV, the drone, and the keys. (i) Robots at their base stations. (ii, iii) Robots searching the apartment. (iv) UGV finds the keys.
V Evaluation Results
The simulation environment features an integrated chat UI through which the human operator (Danny) interacts with the robots. As shown in Fig. 4(1-5), the interface comprises multiple components: a central chat window, visualization widgets for the robots’ language interpretations and visual perceptions (VMRs), natural language representations of robot cognition, and structured displays of their goals and plans. This comprehensive design offers transparency into the robots’ decision-making processes and cognitive operations. The analysis below refers to specific labeled elements in Fig. 4. Please view the accompanying video for a better understanding. Additional information is also available on the HARMONIC project page.  [^1] https://sanjayovs.github.io/HARMONIC2025IROS/
V-1 Task Initiation
@@
The scenario begins with Danny initiating the task by sending the robots a message (M1). This input triggers the team leader (UGV) to place a COLLABORATIVE-ACTIVITY on its agenda (Fig. 4.3) and launch the SEARCH-FOR-LOST-OBJECT plan. Concurrently, the drone awaits instructions.
V-2 Information Gathering
@@
The UGV proceeds to verify preconditions for the SEARCH-FOR-LOST-OBJECT plan (Fig. 4.3). Although the object type is already known, information about its features and last-seen location would be useful. The UGV queries Danny for this information (M2/M4), and receives responses (M3/M5). The latter response prompts the system to prioritize the entry-way in the search sequence (ii).
V-3 Search Execution
@@
With preconditions met, the UGV instructs the drone to initiate the search (M6), and both robots begin exploring their assigned areas as shown in (ii-iv). The individual tactical modules on the robots control the search process using a waypoint strategy, while the strategic (cognitive) module maintains awareness of area existence without directly guiding robot navigation. This approach allows for efficient local path planning while preserving high-level planning in the strategic layer.
V-4 Communication and Coordination
@@
Throughout the search, robots report their findings to each other and the human (M7-8, ii-iv). When a robot fails to locate the keys in a searched area, it communicates this to its partner.
V-5 Task Completion
@@
The VMRs widget for UGV (Fig. 4.2) displays the object detection results that the strategic layer processes from sensing frames communicated by the tactical layer. During the search execution, the strategic module continuously analyzes the sensor data frames, grounding these VMRs against the instance of the KEY object stored in its episodic memory. When the keys are found (iv) – the features match, the search is halted by either of the robots (leader in this case), informs the team (M8), and reports to Danny (M9). Notably, the UGV uses different language constructs when communicating with the drone versus Danny in (M9), demonstrating the cognitive agent’s ability to generate context-appropriate language.
VI Conclusion and Future Work
@@
This work presents a step towards the development of cognitive robots capable of effectively collaborating in heterogeneous human-robot teams. The HARMONIC framework, with its dual-layer architecture that combines strategic cognitive processing and tactical control, allows robots to exhibit metacognitive abilities, including team task understanding, episodic memory utilization, and multimodal perception interpretation. By integrating natural language communication and reasoning capabilities, our approach facilitates complex team interactions, enhances explainability, and builds trust, the key factors in human-robot collaboration.
@@
Our demonstration system of a simulated search task involving two robots and a human teammate illustrates the potential for HARMONIC to be applied to real-world applications. Future research will focus on expanding the range of tasks and scenarios in which HARMONIC agents can successfully operate; enhancing the robots’ learning capabilities; conducting real-world trials using physical robots to validate the framework’s effectiveness in diverse operational contexts; developing collaborative online planning strategies; and enhancing real-time adaptability in dynamic environments.